OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

) SA = 2r(pi) + (pi)r^2 3000 = 2r(pi) + (pi)r^2 (pi)r^2 + 2(pi)r - 3000 = 0 using quadratic formula: r = [-2(pi) +/- SQR((2(pi))^2 - 4(pi)(-3000))]/2(pi) r = [-2(pi) +/- SQR((2(pi))^2 + 12,000(pi))/2(pi) using pi = 3.14: r = [-6.28 +/- SQR(39.4 + 37,680)/6.28 r = [-6.28 +/- 194]/6.28 r = 29.9 or -31.9 using the positive root: r = 29.9cm check: 3000 = 2(29.9)(3.14) + (3.14)(29.9^2) 3000 = 187.8 + 2807.2 3000 = 2995 (very close given the rounding of pi)

OpenStudy (anonymous):

You can take the way I typed it and make it shorter.

OpenStudy (anonymous):

Umm i'll try ?